# TinyGraphEstimator: Adapting Lightweight Language Models for Graph Structure Inference

### v3

@@ 2 RELATED WORK
The task of predicting graph-theoretic properties has traditionally been addressed using specialized graph learning architectures such as Graph Neural Networks (GNNs) [Kipf and Welling, 2016, Hamilton et al., 2017, Wu et al., 2019]. These models operate directly on adjacency structures to learn node embeddings and global representations that capture connectivity and topology. Numerous studies have explored the prediction of structural metrics including clustering coefficients, path lengths, and degree distributions from learned embeddings [You et al., 2018, Dwivedi et al., 2022]. However, while GNNs provide strong inductive biases for graph tasks, they require task-specific supervision and often lack general reasoning ability outside the training distribution. Our work diverges from these approaches by treating the inference of graph properties as a language-based reasoning problem, where information is expressed textually and interpreted by small language models.
@@
Recent advances in large language models (LLMs) have demonstrated impressive performance in tasks requiring symbolic reasoning, numerical computation, and pattern generalization [Brown et al., 2020, Bubeck et al., 2023]. Studies have shown that transformer-based architectures can internalize relational structures [Wang et al., 2025, Schmitt et al., 2021, Ying et al., 2021] and perform graph-related reasoning when provided with suitable prompts or structured input [Huang et al., 2024, Guo et al., 2023, Guo et al., 2025]. Nevertheless, most prior work focuses on large-scale models with hundreds of billions of parameters, whose reasoning capabilities emerge from scale rather than explicit adaptation. In contrast, we investigate whether small language models - under 4B parameters - can infer precise graph-theoretic quantities. This shifts the focus from emergent reasoning in massive models to lightweight, interpretable reasoning in compact architectures.
@@
A growing body of work explores how language models can represent or reason about graphs through textual or hybrid encodings. These approaches encode adjacency information as token sequences, enabling LLMs to reconstruct or classify graph structures [Tang et al., 2024, Chen et al., 2025, Podstawski, 2025]. Other works have investigated relational reasoning benchmarks [Hu et al., 2025, Agrawal et al., 2024], showing that LLMs can approximate certain structural statistics with appropriate context. Our approach builds on this foundation but emphasizes the systematic evaluation of structural parameter inference - quantitative graph descriptors such as density, transitivity, and global efficiency - using small models specifically adapted for this purpose.
@@
Fine-tuning language models for specialized reasoning tasks poses significant computational and memory challenges. Parameter-efficient adaptation techniques such as Low-Rank Adaptation (LoRA) [Hu et al., 2021], prefix tuning [Li and Liang, 2021], and adapter-based methods [Houlsby et al., 2019] have emerged as scalable alternatives that modify only a small subset of parameters while preserving base model capabilities. These techniques have been successfully applied across a range of domains including code understanding [Han et al., 2024], reasoning [Bi et al., 2024], and domain adaptation [Lu et al., 2024]. In our study, we employ LoRA to specialize small language models for structural graph inference, demonstrating consistent performance improvements across all measured graph parameters. This confirms that lightweight fine-tuning can effectively enhance reasoning abilities even in compact architectures.
@@ 3.1 Models Selection
To investigate the ability of small language models to infer structural properties of graphs, we selected a group of compact yet high-performing models that combine efficient deployment with strong reasoning capacity. Our goal was to identify models that are powerful enough to capture graph-theoretic patterns while remaining computationally accessible for systematic experimentation. Model selection was guided by comparative performance metrics reported on the LLM-Stats leaderboard [LLM Stats, 2024], focusing on models with fewer than 4B parameters that demonstrate competitive reasoning accuracy across diverse benchmarks.
The final selection includes Qwen-2.5-3B [Yang et al., 2025], Llama-3.2-3B [Grattafiori et al., 2024], and Phi-4-mini (4B) [Abdin et al., 2024], all used in their Instruct variants. These models were chosen because they consistently rank among the strongest sub-4B architectures in reasoning and comprehension tasks while maintaining full open availability for reproducible research. Their extended context capacity allows complete representations of graphs to be processed in a single inference window. Moreover, their compact architecture enables efficient fine-tuning and evaluation on standard hardware, making them ideal candidates for exploring how small-scale language models internalize and generalize graph-structural information.
For reference, we also evaluate two state-of-the-art LLMs - GPT-4.1 [OpenAI, 2024, Achiam et al., 2024] and DeepSeek-V3.2-Exp (Non-Thinking Mode) [DeepSeek, 2025, Liu et al., 2025] - to benchmark our compact models against the current frontier of LLM capability. While these systems exhibit strong general reasoning and broad-domain adaptability, our fine-tuned small models achieve superior accuracy on graph-structural inference, highlighting the efficiency and specialization benefits of lightweight adaptation.
@@ 3.2 Dataset Construction
To the best of our knowledge, no existing dataset directly supports the task of mapping raw graph structures to such a wide range of quantitative parameters. Therefore, we constructed a balanced synthetic dataset of connected graphs generated from three canonical random graph models: Erdős–Rényi (ER), Barabási–Albert (BA), and Watts–Strogatz (WS). The ER model produces graphs where each possible edge occurs independently with a fixed probability, capturing purely random connectivity [Erdös and Rényi, 1959]. The BA model generates scale-free networks through preferential attachment, emphasizing hub formation and heavy-tailed degree distributions [Barabási and Albert, 1999]. The WS model interpolates between regular lattices and random graphs, combining high clustering with short path lengths [Watts and Strogatz, 1998].
@@
All generated graphs, along with their metadata and manifest files, constitute the TinyGraphEstimator dataset, which is one of the key contributions of this work. It is publicly available to support reproducibility and further research on graph-structured reasoning in language models. The dataset provides a balanced and well-documented collection of connected graphs spanning diverse generative models, enabling systematic evaluation of inference performance across structural regimes.
@@ 3.3 Training Setup
We fine-tune each base model with parameter-efficient supervised instruction tuning using the TRL SFTTrainer [von Werra et al., 2020] and LoRA adapters. Training data consist of paired promptcompletion examples: the prompt includes the normalized graph (edge list), the fixed target schema, and concise instructions; the completion is the gold JSON with expected fields. To avoid learning spurious prompt tokens and to sharpen output formatting, we use completion-only loss. LoRA adapters are applied to attention and MLP projection modules with rank r=32r{=}32, $\alpha{=}32$, and dropout $0.05$. We train with sequence length $2048$, batch size $2$ and gradient accumulation $8$, cosine learning-rate schedule with warmup ratio $0.03$, learning rate $2\!\times\!10^{-4}$, and ten epochs over the training set.
@@ 3.4 Validation Protocol
We evaluate model outputs with a schema- and tolerance-aware validation pipeline that mirrors the training prompt format while remaining model-agnostic. For each sample, the script loads the graph data and gold metadata and normalizes it to a canonical textual form. A structured prompt then instructs the model to return only a JSON object matching a fixed schema of given fields (density, degree statistics, triangles, clustering, transitivity, path lengths, diameter, chromatic number, global efficiency). Inference is run on a base model or a base model augmented with LoRA adapters (if provided), with JSON-constrained decoding via lm-format-enforcer [Gat, 2024] to prevent malformed outputs. The first JSON object in the response is extracted and strictly checked against the schema.
@@ 4 RESULTS
The fine-tuned small models maintain stable accuracy across diverse graph types - from sparse to dense topologies - demonstrating strong generalization and resilience to variations in degree distribution and connectivity. Overall, LoRA-based adaptation enables small language models to reason precisely and interpretably over graph-structured data, establishing them as efficient, high-precision estimators of graph-theoretic properties.
@@ 5 DISCUSSION
The results demonstrate that compact transformer models, when fine-tuned with parameter-efficient techniques such as LoRA, are capable of accurate and interpretable reasoning over graph-structured data. The observed improvements across all graph parameters confirm that small language models can internalize both local and global structural dependencies without requiring explicit graph neural architectures. This finding has several broader implications.
First, it suggests that structural reasoning is not an emergent property of scale alone, but can instead be induced through targeted adaptation. This challenges the assumption that only very large language models possess strong symbolic or numerical inference abilities, and indicates that efficiency-oriented architectures can deliver competitive performance under proper supervision. Second, the TinyGraphEstimator dataset and methodology provide a controlled framework for evaluating how language models interpret formal structures, which could extend beyond graphs to domains such as program analysis, knowledge graphs, or symbolic mathematics. Finally, the results point toward practical opportunities for deploying small LMs in low-resource or embedded environments, where traditional large-scale LLMs are computationally infeasible.
In essence, the strong predictive performance and generalization of fine-tuned models highlight a promising direction for bridging discrete reasoning and natural language modeling. By demonstrating that small transformers can perform accurate structural inference, this work lays the groundwork for future exploration of hybrid symbolic–neural systems and for extending lightweight reasoning models to real-world relational data.
Beyond the presented benchmark, the proposed approach has practical implications for the development of lightweight reasoning agents and embedded AI systems. By enabling small language models to infer and interpret graph-structural properties, TinyGraphEstimator can support tasks such as knowledge graph analysis, network diagnostics, and symbolic planning under resource constraints. These capabilities highlight the potential of compact, fine-tuned models as efficient reasoning modules within larger autonomous or decision-support architectures.
@@ 7 CONCLUSION
This manuscript acknowledges the use of ChatGPT [OpenAI, 2022], powered by the GPT-5 language model developed by OpenAI, to improve language clarity, refine sentence structure, and enhance overall writing precision.

### v4

@@ 2 RELATED WORK
The task of predicting graph-theoretic properties has traditionally been addressed using specialized graph learning architectures such as Graph Neural Networks (GNNs) [kipf2017semi, hamilton2017inductive, wu2020comprehensive]. These models operate directly on adjacency structures to learn node embeddings and global representations that capture connectivity and topology. Numerous studies have explored the prediction of structural metrics including clustering coefficients, path lengths, and degree distributions from learned embeddings [you2018graph, dwivedi2020benchmarking]. However, while GNNs provide strong inductive biases for graph tasks, they often lack general reasoning ability outside the training distribution. Our work diverges from these approaches by treating the inference of graph properties as a language-based reasoning problem, where information is expressed textually and interpreted by small language models.
@@
Recent advances in large language models (LLMs) have demonstrated impressive performance in tasks requiring symbolic reasoning, numerical computation, and pattern generalization [brown2020language, bubeck2023sparks]. Studies have shown that transformer-based architectures can internalize relational structures [wang2025reasoninglargelanguagemodels, schmitt2021modelinggraphstructurerelative, ying2021transformersreallyperformbad] and perform graph-related reasoning when provided with suitable prompts or structured input [huang2024llmseffectivelyleveragegraph, guo2023gpt4graphlargelanguagemodels, guo2025grapheditlargelanguagemodels]. Nevertheless, most prior work focuses on large-scale models with hundreds of billions of parameters, whose reasoning capabilities emerge from scale rather than explicit adaptation. In contrast, we investigate whether small language models - under 4B parameters - can infer precise graph-theoretic quantities. This shifts the focus from emergent reasoning in massive models to lightweight, interpretable reasoning in compact architectures.
@@
A growing body of work explores how language models can represent or reason about graphs through textual or hybrid encodings. These approaches encode adjacency information as token sequences, enabling LLMs to reconstruct or classify graph structures [tang2024graphgptgraphinstructiontuning, chen2025graphgenerativepretrainedtransformer, podstawski2025applyingtextembeddingmodels]. Other works have investigated relational reasoning benchmarks [hu2025rethinkingbenchmarkinglargelanguage, agrawal2024llmsperformstructuredgraph], showing that LLMs can approximate certain structural statistics with appropriate context. Our approach builds on this foundation but emphasizes the systematic evaluation of structural parameter inference - quantitative graph descriptors such as density, transitivity, and global efficiency - using small models specifically adapted for this purpose.
@@
Fine-tuning language models for specialized reasoning tasks poses significant computational and memory challenges. Parameter-efficient adaptation techniques such as Low-Rank Adaptation (LoRA) [hu2021loralowrankadaptationlarge], prefix tuning [li2021prefixtuningoptimizingcontinuousprompts], and adapter-based methods [houlsby2019parameterefficienttransferlearningnlp] have emerged as scalable alternatives that modify only a small subset of parameters while preserving base model capabilities. These techniques have been successfully applied across a range of domains including code understanding [han2024parameterefficientfinetuninglargemodels], reasoning [bi2024enhancingreasoningcapabilitiessmall], and domain adaptation [lu2024finetuninglargelanguagemodels]. In our study, we employ LoRA to specialize small language models for structural graph inference, demonstrating consistent performance improvements across all measured graph parameters. This confirms that lightweight fine-tuning can effectively enhance reasoning abilities even in compact architectures.
@@ 3.1 Models Selection
To investigate the ability of small language models to infer structural properties of graphs, we selected a group of compact yet high-performing models that combine efficient deployment with strong reasoning capacity. Our goal was to identify models that are powerful enough to capture graph-theoretic patterns while remaining computationally accessible for systematic experimentation. Model selection was guided by comparative performance metrics reported on the LLM-Stats leaderboard [llmstats_github_2024], focusing on models with fewer than 4B parameters that demonstrate competitive reasoning accuracy across diverse benchmarks.
The final selection includes Qwen-2.5-3B [qwen2025qwen25technicalreport], Llama-3.2-3B [grattafiori2024llama3herdmodels], and Phi-4-mini (4B) [abdin2024phi4technicalreport], all used in their Instruct variants. These models were chosen because they consistently rank among the strongest sub-4B architectures in reasoning and comprehension tasks while maintaining full open availability for reproducible research. Their extended context capacity allows complete representations of graphs to be processed in a single inference window. Moreover, their compact architecture enables efficient fine-tuning and evaluation on standard hardware, making them ideal candidates for exploring how small-scale language models internalize and generalize graph-structural information.
For reference, we also evaluate two state-of-the-art LLMs - GPT-4.1 [openai_gpt41_2024, openai2024gpt4technicalreport] and DeepSeek-V3.2-Exp (Non-Thinking Mode) [deepseek_docs_2025, deepseekai2025deepseekv3technicalreport] - to benchmark our compact models against the current frontier of LLM capability. While these systems exhibit strong general reasoning and broad-domain adaptability, our fine-tuned small models achieve superior accuracy on graph-structural inference, highlighting the efficiency and specialization benefits of lightweight adaptation.
@@ 3.2 Dataset Construction
To the best of our knowledge, no existing dataset directly supports the task of mapping raw graph structures to such a wide range of quantitative parameters. Therefore, we constructed a balanced synthetic dataset of connected graphs generated from three canonical random graph models: Erdős–Rényi (ER), Barabási–Albert (BA), and Watts–Strogatz (WS). The ER model produces graphs where each possible edge occurs independently with a fixed probability, capturing purely random connectivity [erdos59a]. The BA model generates scale-free networks through preferential attachment, emphasizing hub formation and heavy-tailed degree distributions [barbasi_1999]. The WS model interpolates between regular lattices and random graphs, combining high clustering with short path lengths [watts_collective_1998].
@@
All generated graphs, along with their metadata and manifest files, constitute the TinyGraphEstimator dataset, one of the key contributions of this work. It is publicly accessible [^1] https://doi.org/10.5281/zenodo.17387934 to support reproducibility and further research on graph-structured reasoning in language models.
@@ 3.3 Training Setup
We fine-tune each base model with parameter-efficient supervised instruction tuning using the TRL SFTTrainer [vonwerra2022trl] and LoRA adapters. Training data consist of paired promptcompletion examples: the prompt includes the normalized graph (edge list), the fixed target schema, and concise instructions; the completion is the gold JSON with expected fields. To avoid learning spurious prompt tokens and to sharpen output formatting, we use completion-only loss. LoRA adapters are applied to attention and MLP projection modules with rank r=32r{=}32, $\alpha{=}32$, and dropout $0.05$. We train with sequence length $2048$, batch size $2$ and gradient accumulation $8$, cosine learning-rate schedule with warmup ratio $0.03$, learning rate $2\!\times\!10^{-4}$, and ten epochs over the training set.
@@ 3.4 Validation Protocol
We evaluate model outputs with a schema- and tolerance-aware validation pipeline that mirrors the training prompt format while remaining model-agnostic. For each sample, the script loads the graph data and gold metadata and normalizes it to a canonical textual form. A structured prompt then instructs the model to return only a JSON object matching a fixed schema of given fields (density, degree statistics, triangles, clustering, transitivity, path lengths, diameter, chromatic number, global efficiency). Inference is run on a base model or a base model augmented with LoRA adapters (if provided), with JSON-constrained decoding via lm-format-enforcer [noamgat2023lmformatenforcer] to prevent malformed outputs. The first JSON object in the response is extracted and strictly checked against the schema.
@@ 4 RESULTS
The fine-tuned small models maintain stable accuracy across diverse graph types - from sparse to dense topologies - demonstrating strong generalization and resilience to variations in degree distribution and connectivity. These outcomes highlight the effectiveness of LoRA-based adaptation in enhancing the structural inference capabilities of compact language models.
@@ 5 DISCUSSION
The results highlight how targeted, parameter-efficient adaptation enables small transformer models to acquire reliable structural inference skills from textual graph representations. Consistent improvements across all evaluated graph parameters confirm that small language models are capable of internalizing both local and global structural dependencies without requiring explicit graph neural architectures.
The ability of compact models to infer structural patterns can be attributed to several complementary mechanisms. First, the self-attention mechanism implicitly constructs relational dependencies between tokens, enabling recovery of adjacency-like patterns from serialized edge lists. Second, weight sharing and low-rank adaptation promote compression of relational rules, supporting the discovery of generalizable structural regularities rather than memorization. Third, pretrained linguistic priors - such as token co-occurrence, dependency tracking, and approximate counting - introduce transferable inductive biases that facilitate reasoning over graph-like inputs. Finally, the hierarchical organization of attention layers allows aggregation from local edge information to global structural descriptors, enabling accurate estimation of higher-order metrics such as clustering, transitivity, and global efficiency. Together, these mechanisms explain how small transformers, despite limited capacity and the absence of explicit graph inductive biases, can learn and generalize structural patterns directly from textual graph representations.
These findings have implications for both theory and methodology. They suggest that structural reasoning is not solely an emergent property of model scale but can also be encouraged through targeted adaptation, indicating that compact, efficiency-oriented architectures are capable of meaningful reasoning performance when appropriately fine-tuned. Building on this insight, the TinyGraphEstimator framework provides a controlled and reproducible environment for investigating such behaviors within the domain of graph reasoning, offering a principled way to examine how small language models interpret and generalize structural patterns. Although this study focuses on graph-structured data, the same methodological principles could extend to other structured domains in which relational dependencies are fundamental.
Beyond conceptual insights, these results suggest practical opportunities for deploying small language models in resource-constrained or embedded environments. Fine-tuned compact transformers can serve as efficient reasoning components within larger autonomous or decision-support systems, supporting applications such as knowledge graph analysis, network diagnostics, and symbolic planning. By enabling small models to infer and interpret graph-structural properties effectively, TinyGraphEstimator highlights a pathway toward lightweight reasoning agents that combine interpretability, adaptability, and scalability. Overall, the demonstrated predictive performance and generalization ability of fine-tuned small models help bridge the gap between discrete symbolic reasoning and natural language modeling, establishing compact transformers as viable tools for structured inference across both research and applied contexts.
@@ 7 CONCLUSION
This manuscript acknowledges the use of ChatGPT [chatgpt], powered by the GPT-5 language model developed by OpenAI, to improve language clarity, refine sentence structure, and enhance overall writing precision.